# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ 5.2 Local dependence on parameters of the slow solution
Let $t_{0}>0$, $\boldsymbol{q}\in\mathbb{R}^{m}$, and $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ denote a pair of primal and dual solutions to BPDN at hyperparameter $t_{0}$ and data $\boldsymbol{b}$. In addition, let
$\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{v}+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})}\right\|_{2}^{2}$ (35)
subject $\text{subject to}\ \begin{cases}\boldsymbol{v}_{j}&\geqslant 0\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})=0$},\\ \boldsymbol{v}_{j}&=0\ \text{if $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$},\end{cases}$
@@
$\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q}),$ (36)
$C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq$
$\quad\inf_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left(\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle\right)}-\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle}\right\},$
$T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\frac{t_{0}}{1+t_{0}C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})},$
$T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})\coloneqq t_{0}\left(1-\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}_{j}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|}\right),$
$t_{1}\coloneqq\max\left(T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s}),T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\right).$
Then, we have the following:
  • (i)
    There holds $0\leqslant t_{1}<t_{0}$, and the primal and dual solutions to basis pursuit denoising at hyperparameter $t\in[t_{1},t_{0}]$ and data $\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}$ are given by
    $\boldsymbol{x}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ $=\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),$ (37)
    $\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ =$=\begin{cases}&\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\quad\text{if $t_{1}(t_{0},\boldsymbol{b},\boldsymbol{q})>0$},\\ &\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\quad\text{otherwise.}\end{cases}$
  • (ii)
    (Monotonicity) Assume $\boldsymbol{q}=-\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\boldsymbol{w}/t_{0}$ for some nonnegative $\boldsymbol{w}\in\mathbb{R}^{n}$. Then
    $\begin{gathered}V(\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})\leqslant V(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \left\|{\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}\right\|_{2}\leqslant\left\|{\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})}\right\|_{2}.\end{gathered}$ (38)
    If also $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\neq\boldsymbol{0}$, then both inequalities in (38) are strict.
@@
See Appendix B.4 for the proof. ∎
@@
The numbers $T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})$ and $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})$ identify potential “kinks” in the piecewise linear dependence of the primal and dual solutions on $t$. If $t_{1}=T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})>0$, then there is at least one index $j\in\{\{1,\dots,n\}:\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})=0\}$ that joins the set $\mathcal{E}(\boldsymbol{p}^{s}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})\boldsymbol{q}))$, and this index is new if $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$. If $t_{1}=T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})>0$, then there is at least one index $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ such that $\boldsymbol{x}^{s}_{j}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})\boldsymbol{q})=\boldsymbol{0}$. This index may leave the equicorrelation set $\mathcal{E}(\boldsymbol{p}^{s}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})))$ beyond $t>t_{1}$.
5.3 Non-uniqueness in the local dependence on parameters
The local continuation of $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ depends on the solution $\boldsymbol{v}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ and residual vector $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ of the NNLS problem (35). Since the residual vector is always unique, both the number $T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})$ and the continuation of $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$ are also unique. The NNLS solution, however, is generally not unique. This leads to complications because both the number $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})$ and continuation of $\boldsymbol{x}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ beyond $t<t_{0}$ may not be unique [^2] This problem was pointed out earlier by Bringmann et al. [11] in a different context, where the authors constructed a homotopy algorithm directly from the KKT conditions of BPDN. Their NNLS problem (Equation 3.2. on page 6) is equivalent to (35) with $\boldsymbol{q}=\boldsymbol{0}$.. This is the well-known problem of non-uniqueness of primal solutions to BPDN [63].
@@ 5.3 Non-uniqueness in the local dependence on parameters
$\hat{\boldsymbol{v}}^{\text{LSQ}}=t_{0}(\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}\bm{A}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))})^{-1}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})$
@@
5.4 Global dependence on parameters of the slow solution
From a dynamical systems perspective, Lemma 5.1 and Proposition 5.1 describe the local continuation of the asymptotic limit of the slow system (12) in terms of the hyperparameter and data. This continuation is local because it holds on an interval $t\in[t_{1},t_{0}]$. As discussed in Section 5.3, the continuation is not unique if the solution to the NNLS problem (35) is not unique. Nevertheless, if $t_{1}\neq 0$, we can apply Proposition 5.1 again to extend the local continuation to an interval $[t_{2},t_{0}]$ with $0\leqslant t_{2}<t_{1}$.
We now apply Proposition 5.1 repeatedly with $\boldsymbol{q}=\boldsymbol{0}$ throughout, starting from $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$, and $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. Doing so $k$ times yields a global continuation of the asymptotic limit of the slow system (12), hence a global continuation of the pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ on $t\in[t^{(k)},t^{(0)}]$.
As discussed in Section 5.3, the continuation in the primal solution $\boldsymbol{x}^{s}(t,\boldsymbol{b})$ is not unique. Choosing at each $k$ the NNLS solution to problem (35) with least $\ell_{2}$ norm, we obtain Algorithm 2 below. It is nearly identical to the generalized homotopy algorithm proposed in Bringmann et al. [11], with the exception of the breakpoints, which here are simpler and more explicit. Algorithm 2 yields a solution path of BPDN and converges in finite time.
@@ 5.4 Global dependence on parameters of the slow solution
Output : A finite sequence $\left\{t^{(k)},\boldsymbol{x}^{(k)},\boldsymbol{p}^{(k)}\right\}_{k=0}^{K}$ specifying a solution path to the basis pursuit denoising problem.
1 Set $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$ and $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$;
@@
7       Compute the numbers $T_{-}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0},\hat{\boldsymbol{v}}^{(k-1)})$ and $T_{+}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0})$ from (36);
8       Update $t^{(k)}=\max(T_{-}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0},\hat{\boldsymbol{v}}^{(k-1)}),T_{+}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0}))$;
@@ 15             Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+(1/t^{(k)}-1/t^{(k-1)})\boldsymbol{\xi}^{(k-1)}$;
Algorithm 2 Homotopy algorithm for computing the primal/dual solution paths to the basis pursuit denoising problem.
@@
Correctness follows immediately from Proposition 5.1 and the discussion in Section 5.3. The proof of convergence in finite time is identical to that of Theorem 4.2 in Bringmann et al. [11] and is omitted. ∎

### v2

@@ 5.1 Local dependence on hyperparameter of the slow solution
Let $t_{0}>0$ and $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ denote a pair of primal and dual lasso solutions at hyperparameter $t_{0}$ and data $\boldsymbol{b}$. In addition, let
$\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{v}+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\right\|_{2}^{2}$ (36)
[…]
@@
$\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})\coloneqq\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})),$ (37)
$C^{s}(t_{0},\boldsymbol{b})\coloneqq$
$\quad\inf_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left(\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\right)}-\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle}\right\},$
$T_{+}(t_{0},\boldsymbol{b})\coloneqq\frac{t_{0}}{1+t_{0}C^{s}(t_{0},\boldsymbol{b})},\qquad T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})\coloneqq t_{0}\left(1-\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}_{j}^{s}(t_{0},\boldsymbol{b})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b})|}\right),$
$t_{1}\coloneqq\max\left(T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s}),T_{+}(t_{0},\boldsymbol{b})\right).$
Then $0\leqslant t_{1}<t_{0}$ and, for every $t\in[t_{1},t_{0}]$,
$\boldsymbol{x}^{s}(t,\boldsymbol{b})$ $=\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b}),$ (38)
[…]
@@
See Appendix B.3 for the lengthy and technical proof. ∎
@@
The numbers $T_{+}(t_{0},\boldsymbol{b})$ and $T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})$ identify potential “kinks” in the piecewise linear dependence of the primal and dual solutions on $t$. If $t_{1}=T_{+}(t_{0},\boldsymbol{b})>0$, then there is at least one index $j\in\{\{1,\dots,n\}:\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})=0\}$ that joins the set $\mathcal{E}(\boldsymbol{p}^{s}(t_{1},\boldsymbol{b}))$, and this index is new if $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$. If $t_{1}=T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})>0$, then there is at least one index $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ such that $\boldsymbol{x}^{s}_{j}(t_{1},\boldsymbol{b})=\boldsymbol{0}$. This index may leave the equicorrelation set $\mathcal{E}(\boldsymbol{p}^{s}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})))$ beyond $t>t_{1}$.
5.2 Non-uniqueness in the local dependence on hyperparameter
The local continuation of $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ depends on the solution $\boldsymbol{v}^{s}(t_{0},\boldsymbol{b})$ and residual vector $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})$ of the NNLS problem (36). Since the residual vector is unique, the number $T_{+}(t_{0},\boldsymbol{b})$ and the continuation of $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$ are also unique. The NNLS solution, however, is generally not unique. This leads to complications because the number $T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})$ may not be unique. Thus the continuation of $\boldsymbol{x}^{s}(t,\boldsymbol{b})$ beyond $t<t_{0}$ may not be unique. This is not new and is a well-known problem [48, 9].
@@ 5.2 Non-uniqueness in the local dependence on hyperparameter
$\hat{\boldsymbol{v}}^{\text{LSQ}}=t_{0}(\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}\bm{A}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))})^{-1}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$
@@
5.3 Global dependence on hyperparameter of the slow solution
From a dynamical systems perspective, Lemma 5.1 and Proposition 5.1 describe the local continuation of the asymptotic limit of the slow system (12) in terms of the hyperparameter. This continuation is local because it holds on an interval $t\in[t_{1},t_{0}]$, and it is not unique whenever the solution to the NNLS problem (36) is not unique. Nevertheless, if $t_{1}\neq 0$, we can apply Proposition 5.1 again to extend the local continuation to an interval $t_{2},t_{0}]$ with $0\leqslant t_{2}<t_{1}$.
We now apply Proposition 5.1 repeatedly, starting from $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$, and $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. Doing so $k$ times yields a global continuation of the asymptotic limit of the slow system (12), hence a global continuation of the pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ on $t\in[t^{(k)},t^{(0)}]$. As discussed in Section 5.2, the continuation in the primal solution $\boldsymbol{x}^{s}(t,\boldsymbol{b})$ is not unique. Choosing at each $k$ the NNLS solution to problem (36) with least $\ell_{2}$ norm, we obtain Algorithm 2 below. It is nearly identical to the generalized homotopy algorithm proposed by Bringmann et al. [9], with the exception of the breakpoints, which here are simpler and more explicit. Algorithm 2 yields a lasso solution path and converges in finite time.
@@ 5.3 Global dependence on hyperparameter of the slow solution
Output : A finite sequence $\left\{t^{(k)},\boldsymbol{x}^{(k)},\boldsymbol{p}^{(k)}\right\}_{k=0}^{K}$ specifying a solution path to the lasso problem.
1
21exSet $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$ and $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$;
@@
8   Compute the numbers $T_{-}(t^{(k-1)},\boldsymbol{b},\hat{\boldsymbol{v}}^{(k-1)})$ and $T_{+}(t^{(k-1)},\boldsymbol{b})$ from (37);
9   Update $t^{(k)}=\max(T_{-}(t^{(k-1)},\boldsymbol{b},\hat{\boldsymbol{v}}^{(k-1)}),T_{+}(t^{(k-1)},\boldsymbol{b}))$;
@@ 16      Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+(1/t^{(k)}-1/t^{(k-1)})\boldsymbol{\xi}^{(k-1)}$;
Algorithm 2 Homotopy algorithm for computing the primal and dual solution paths to the lasso problem.
@@
Correctness follows from Proposition 5.1 and the discussion in Section 5.2. Convergence in finite time is identical to that of Theorem 4.2 in Bringmann et al. [9] and is omitted. ∎